Required information [The following information applies to the questions displayed below.]

In 2018, Elaine paid $2,840 of tuition and $800 for books for her dependent son to attend State University this past fall as a freshman. Elaine files a joint return with her husband. What is the maximum American opportunity credit that Elaine can claim for the tuition payment and books in each of the following alternative situations?

b. Elaine’s AGI is $172,000. (Round your intermediate calculations to the nearest whole dollar amount.)
Answer #1

ANSWER

Qualifying education expense = 2,840 -800 = $2,040

American opportunity credit is 100% of the first $2,000 and 25% of the next 2000 of qualifying education expenses.

The AGI of Elaine is less than the phase-out threshold of $160,000.

So, the full amount of eligible credit can be claimed.

American opportunity credit is

= 2000+(2,040-2000)*25%

= $2,010.

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************